The Baldwin company wants to decrease its plant utilization for Buzz by 15%. How many units would need to be produced next year to meet this production goal? Ignore impact of accounts payable on plant utilization. 1) 1,211 2) 1,029 3) 1,661 4) 1,436The Baldwin company wants to decrease its plant ut
Answer #1
Concepts and reason

Capacity Utilization: It is the extent to which a company might utilize its installed production capacity. A company might alter its production capacity based on the demand and supply of raw material.

Fundamentals

Direct materials: These are the materials which are used in the Production process which constitutes an integral portion of the final product, the cost of which is identifiable and traceable directly to it. The cost of direct material can be directly chargeable to the final product as compared to the indirect material.

Direct labor: Workers or employees directly involved in the manufacture of goods and services are called the direct labor. Direct labor can be directly traceable to a specific product or cost center. In other words, the cost of the workers which is directly involved in the transformation of the raw material into the finished product.

Contribution Margin: It represents the excess of sales over its variable cost. It judges whether the company is able to cover its variable cost and contributes towards the fixed cost.

It is computed by subtracting the variable cost from sales revenue.

Production budget: Production budget is prepared according to plan production in units which must be made by manufacturers in a specific time period. This budget is predicted by the sales target.

Calculate the new plant utilization rate using the equation as shown below.

NewPlantutilizationrate=ExistingPlantutilizationDecreaseinplantutilization=96%15%=81%\begin{array}{c}\\{\rm{New Plant utilization rate}} = {\rm{Existing Plant utilization}} - {\rm{Decrease in plant utilization}}\\\\ = {\rm{96\% }} - {\rm{15\% }}\\\\ = {\rm{81\% }}\\\end{array}

The new plant utilization rate is 81%.

Calculate the production target for the new capacity rate using the equation as shown below.

Productiontarget=Totalplantcapacity×Capacityutilizationrate=1,500units×81%=1,500units×81100=1,215units\begin{array}{c}\\{\rm{Production target}} = {\rm{Total plant capacity}} \times {\rm{Capacity utilization rate}}\\\\ = {\rm{1,500 units}} \times {\rm{81\% }}\\\\ = {\rm{1,500 units}} \times \frac{{81}}{{100}}\\\\ = 1,21{\rm{5 units}}\\\end{array}

The production target for the new capacity rate is 1,215 units

Nearest option available is 1,211. So, the correct answer is 1,211 units.

Ans:

The units required to be produced to meet the target at 81% capacity is 1,211 units.